In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?
Answer by rapaljer(4671)

Let x = number of votes of the losing candidate
x+220 = number of votes of the winning candidate

x+x+220=810
2x + 220 = 810
2x = 810-220
2x = 590
x = 295 Loser
x+220 = 515 Winner

Check: 295 + 220 = 515

R^2